The area of a metal is given as 37 ft2.  What is its area in  a) in2,     b) cm2, and c) m2?

 



1
Expert's answer
2021-08-29T16:56:29-0400

(a)

"37\ ft^2\cdot(\dfrac{12\ in}{1\ ft})^2=5328\ in^2."

(b)

"37\ ft^2\cdot(\dfrac{30.48\ cm}{1\ ft})^2=34374\ cm^2."

(c)

"37\ ft^2\cdot(\dfrac{0.3048\ m}{1\ ft})^2=3.44\ m^2."